Nov 6, 10:08 UTC The fiber has been changed and the switchover done. We are monitoring the situation. Nov 6, 08:56 UTC We are getting reports about connection problems in the data center #2. While we do not see outages there are recurring spikes in latency, we see between the data center MUC1 and MUC2 interconnected by dark fiber. We have already switched over to another switch and changed transceivers, but the problem seems to persist. Therefore, we will now change the fiber itself connecting the main switches between our data centers. This switchover could potentially cause up to 90 seconds of downtime for all servers and services located in the data center MUC2. The work will be carried out at approx. 10:00 UTC (approx. 1 hour from this announcement).